# ShadowMaskFormer: Mask Augmented Patch Embedding for Shadow Removal

### v2

@@ 6 Ablation Studies
Ablation Study on MAPE. As summarized in Table 4, we first present the performance of the regular path embedding [51], which is the de-facto choice of image preprocessing adopted by most existing works that are based on transformers. Evidently, our mask augmented patch embedding (MAPE) leads to significantly better shadow removal performance measured by RMSE. Compared to the conventional patch embedding, our MAPE can obviously utilize the regional information provided by the shadow mask effectively and convey this information to the entire model. Then, we present the relative improvements from the proposed binarization scheme by replacing it with the regular 0/1 binary mask (i.e., ), where the effectiveness of can be observed. This experimental result demonstrates that the combination of the two shadow masks designed in our MAPE is significantly more effective in shadow removal compared to using only a single, common 0/1 mask (). In other words, the success of MAPE relies on the reliable combined use of the two masks. On the contrary, applying shadow masks directly within patch embedding can lead to performance degradation in shadow regions, as shown in Table 4.
Method S NS All RMSE RMSE RMSE Original PE 7.59 (+1.5) 4.91 (+1.1) 5.35 (+1.1) MAPE 6.08 (+0.0) 3.86 (+0.0) 4.23 (+0.0) in MAPE 8.03 (+2.0) 4.49 (+0.6) 4.95 (+0.7)
@@
Model Size Params S NS All (M) RMSE RMSE RMSE Small 1.94 6.61 3.93 4.37 Middle 2.28 6.08 3.86 4.23 Large 2.52 6.27 3.80 4.20

### v3

@@ 6 Ablation Studies
Ablation Study on MAPE. As summarized in Table 4, we first present the performance of the regular path embedding [53], which is the de-facto choice of image preprocessing adopted by most existing works that are based on transformers. Evidently, our proposed mask augmented patch embedding leads to significantly better shadow removal performance measured by MAE. Compared to the conventional patch embedding, our MAPE can obviously utilize the regional information provided by the shadow mask effectively and convey this information to the entire model. Then, we present the relative improvements from the proposed binarization scheme by replacing it with the regular 0/1 binary mask (i.e., ), where the effectiveness of can be observed. This experimental result demonstrates that the combination of the two shadow masks designed in our MAPE is significantly more effective in shadow removal compared to using only a single, common 0/1 mask (). In other words, the success of MAPE relies on the reliable combined use of the two masks. On the contrary, applying shadow masks directly within patch embedding can lead to performance degradation in shadow regions, as shown in Table 4.
Method S NS All MAE MAE MAE Original PE 7.59 (+1.5) 4.91 (+1.1) 5.35 (+1.1) MAPE 6.08 (+0.0) 3.86 (+0.0) 4.23 (+0.0) in MAPE 8.03 (+2.0) 4.49 (+0.6) 4.95 (+0.7)
@@
Model Size Params S NS All (M) MAE MAE MAE Small 1.94 6.61 3.93 4.37 Middle 2.28 6.08 3.86 4.23 Large 2.52 6.27 3.80 4.20
The effect of the quality of masks. Evaluating the robustness of MAPE to varying-quality shadow masks is crucial. We conducted a series of experiments and analyses to evaluate the robustness of MAPE to varying-quality shadow masks. Results suggest that MAPE maintains strong performance, even with inaccurate or low-quality masks. Specifically, the SRD dataset [5] uses shadow masks predicted by DHAN [10], which are incomplete and inaccurate, as shown in Figure 5. Despite this, our method achieves efficient shadow removal, demonstrating MAPE’s robustness to inaccurate input masks. To further quantify this robustness, we degraded the shadow masks in the ISTD dataset [8] to generate low-quality masks. Specifically, since the accuracy of shadow masks is related to the correctness of shadow boundaries, we first extracted the shadow region contours from the dataset’s shadow masks. We then randomly misclassified contour shadow pixels as non-shadow pixels to generate low-quality masks. As shown in the first row of Figure 6, different low-quality masks are presented, where the differences between low-quality and ISTD dataset masks are visible. The quality of these masks was evaluated using the balance error rate (BER) [54], where lower BER indicates higher accuracy. The results presented in Table 6 indicate that shadow removal performance is closely related to mask quality. When the BER increases, reflecting lower mask quality, the shadow removal performance degrades. However, even when the BER increases sevenfold (from 0.61 to 4.39), the MAE only changes by 11%, demonstrating that MAPE maintains its robustness to varying-quality masks. Finally, visual comparisons of shadow masks with different quality and the corresponding results are provided in Figure 6. We can observe slight contour issues in the shadow regions when BER is 0.61 or 1.82, and more pronounced issues when BER reaches 4.39.
Mask S NS All BER MAE MAE MAE 0.61 7.13 4.19 4.67 1.29 7.36 4.21 4.72 1.82 7.78 4.20 4.78 2.85 8.47 4.22 4.93 3.36 8.87 4.24 5.02 4.39 9.72 4.29 5.21
Generalization Performance of ShadowMaskFormer. To assess the generalization capability of ShadowMaskFormer, we conducted experiments on previously unseen scenarios. Specifically, we selected images from the ISTD dataset where shadows are cast on walls, an extremely common real-world scenario, as the unseen test set, while the remaining images were used for training. We maintained the same experimental setup as described in Section 5.1, ensuring consistency in the data partitioning strategy and evaluation approach. Unlike artificial or synthetic test settings, wall shadows naturally occur in diverse environments with textured surfaces and varying reflectance properties, providing a representative benchmark for evaluating the generalizability of shadow removal models in practical applications. The results, presented in Table 7, show that ShadowMaskFormer refers to the model trained and tested on the ISTD dataset, while ShadowMaskFormer* refers to the generalized results. We observe that ShadowMaskFormer* shows a slight decrease in performance in the shadowed regions, while performance in non-shadow regions improves, with overall image performance remaining consistent. These findings suggest that our method effectively adapts to unseen shadow scenarios, handling novel shadow characteristics while maintaining robust global performance.
Methods S NS All MAE MAE MAE ShadowMaskFormer 6.08 3.86 4.23 ShadowMaskFormer* 6.69 3.55 4.23
Additionally, as shown in Figure 7, the visual results further demonstrate that our method successfully removes shadows even in previously unseen settings. This highlights the strong generalization capability of ShadowMaskFormer, enabling it to handle real-world shadow removal tasks effectively.